# A Unified Theory of Sparse Dictionary Learning in Mechanistic Interpretability: Piecewise Biconvexity and Spurious Minima

### v3

@@ Abstract
Sparse Autoencoders, Dictionary Learning, Interpretability, Neural Networks
@@ 1 Introduction
Despite this empirical success across diverse applications, SDL methods consistently exhibit persistent failure modes: learned features remain polysemantic (Chanin et al., 2025), ”dead neurons” fail to activate on any data samples (Bricken et al., 2023), and ”feature absorption” occurs where one neuron captures specific sub-concepts while another responds to the remaining related concepts (Chanin et al., 2025). Practitioners have developed techniques to address these issues, including neuron resampling (Bricken et al., 2023), auxiliary losses (Gao et al., 2024), and careful hyperparameter tuning, yet these fixes remain ad-hoc engineering solutions without principled justification. Critically, these phenomena persist even with careful training on clean data, suggesting they are not mere implementation artifacts but reflect fundamental structural properties of the SDL optimization problem itself.
@@
In this work, we develop a unified theoretical framework that formalizes SDL as a general optimization problem, encompassing various SDL methods (Bussmann et al., 2024, 2025; Tang et al., 2025b) as special cases. We demonstrate how these diverse methods instantiate our framework through different choices of input-output representation pairs, activation functions, and loss designs. We establish rigorous conditions under which SDL methods provably recover ground-truth interpretable features, characterizing the roles of feature sparsity, latent dimensionality, and activation functions. Through detailed analysis of the optimization landscape, we demonstrate that global minima correspond to correct feature recovery and provide necessary and sufficient conditions for achieving zero loss. We establish the prevalence of spurious partial minima exhibiting un-disentangled polysemanticity, providing novel theoretical explanations for feature absorption (Chanin et al., 2025) and the effectiveness of neuron resampling (Bricken et al., 2023). We design the Linear Representation Bench, a synthetic benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theoretical insights, we propose feature anchoring, a technique applicable to all SDL methods which achieves improved feature recovery by constraining learned features to known anchor directions.
@@ 2 Preliminaries
In this section, we present a unified theoretical framework for Sparse Dictionary Learning (SDL). We begin with the formal definitions of foundational concepts and the linear representation hypothesis. Then we introduce our framework and how various SDL methods instantiate it.
@@ 2.1 Input Distribution and Model Representation
2.2 The Linear Representation Hypothesis
Empirical studies in mechanistic interpretability have observed that neural network representations encode meaningful concepts as linear directions, often in superposition (Marks and Tegmark, 2024; Nanda et al., 2023; Jiang et al., 2024; Park et al., 2025). Following Elhage et al. (2022) and Park et al. (2024), we formalize the hypothesis as follows.
Assumption 2.3 (Linear Representation Hypothesis).
A model representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ satisfies the Linear Representation Hypothesis (LRH) if there exists a feature function $\mathbf{x}:\mathcal{X}\to\mathbb{R}^{n}$ and a feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ such that:
@@ 2.2 The Linear Representation Hypothesis
  4. 4.
    Interpretability: Each component $x_{i}$ of $\mathbf{x}$ corresponds to a human-interpretable concept.
We refer to $\mathbf{x}(s)$ as the ground-truth features, where each $x_{i}(s)$ represents the activation level of concept $i$ for input $s$. $W_{p}$ encodes how these features are represented in $\mathbf{x}_{p}$.
Assumption 2.4 (Representation Assumptions).
Following recent works (Elhage et al., 2022; Park et al., 2024), we assume the following conditions on the model representations satisfying the Linear Representation Hypothesis, which align with practical use cases:
  1. 1.
    $\|W_{p}[:,i]\|_{2}=1\quad\forall i\in[n]$
  2. 2.
    Extreme Sparsity: The sparsity level $S$ approaches 1:
    $S\to 1$
  3. 3.
    Independence: The ground-truth features satisfy: For any input $s\sim\mathcal{D}$, the features $\{x_{i}(s)\}_{i=1}^{n}$ are mutually independent and independent of the sparsity level $S$.
  4. 4.
    Bounded Interference: Define the maximum interference as:
    $M:=\max_{i\neq j}\langle W_{p}[:,i],W_{p}[:,j]\rangle$
where $W_{p}[:,i]$ denotes the $i$-th column of $W_{p}$. In typical superposition scenarios, $M>0$ but remains small (close to zero), characterizing a regime where many interpretable features are compressed into fewer dimensions with bounded mutual interference.
@@ 2.3 General Optimization Framework for SDL
Now we formalize SDL as an optimization problem under the Linear Representation Hypothesis (Assumption 2.3) and the Representation Assumptions (Assumption 2.4).
@@ 2.4 Instantiations: Existing SDL Methods
This property holds for ReLU, JumpReLU, Top-$k$, Batch Top-$k$, and their compositions, the primary sparsity mechanisms in practice. Notably, all these activation functions can be expressed in the form $\sigma_{\text{Jump}}(z;c)$ for some threshold $c\geq 0$: $\sigma_{\text{Jump}}(z;c)_{i}=z_{i}\cdot\mathbf{1}(z_{i}>c)$.
@@ 3.1 Approximate Loss with Feature Reconstruction
Under extreme sparsity, the SDL loss decomposes into independent per-feature reconstruction terms.
@@
Under Assumptions 2.3 and 2.4, define the approximate loss as:
@@
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$. Then
$|\mathcal{L}_{\text{SDL}}-\tilde{\mathcal{L}}_{\text{SDL}}|\leq O((1-S)^{2})$ (12)
This approximation isolates the contribution of each ground-truth feature, making the optimization landscape amenable to convex analysis techniques applied subsequently.
@@ 3.2 SDL Optimization is Piecewise Biconvex
Consider the approximate SDL loss $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$. Define the activation pattern region as $\Omega_{\mathcal{A}}=\{W_{E}\in\mathbb{R}^{n_{q}\times n_{p}}:\forall d\in[n],\mathcal{A}(d)=\{i\in[n_{q}]:\langle w_{E}^{i},w_{p}^{d}\rangle>c\}\}$ where $c\geq 0$ is the activation threshold and $\mathcal{A}(d)\subseteq[n_{q}]$ denotes neurons activated by feature $d$.
Then $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibits bi-convex structure over $\mathbb{R}^{n_{r}\times n_{q}}\times\Omega_{\mathcal{A}}$:
@@
    For any fixed $W_{E}\in\Omega_{\mathcal{A}}$, $W_{D}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{D}$.
@@
    For any fixed $W_{D}\in R^{nr\texttimes nq},W_{E}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{E}$ over $\Omega_{\mathcal{A}}$.
This establishes SDL as a biconvex optimization problem, bridging mechanistic interpretability with classical biconvex optimization theory. This characterization enables both theoretical analysis of the optimization landscape and implementation of specialized biconvex methods for SDL.
@@ 3.3 Characterizing the Global Minimum of SDL
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$.
When $n_{q}\geq n$ and $\sigma$ satisfies $\sigma(z)_{i}\in\{0,z_{i}\}$ for all $i$, the configuration
@@
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=O((1-S)M^{2})$ (15)
where $M$ is the maximum interference.
@@ 3.4 Characterizing Spurious Partial Minima of SDL
Under Assumptions 2.3 and 2.4 with $n\geq 2$ and $n_{q}\geq n$, for any activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ that is realizable, polysemantic, and forms a partition of $[n]$, there exists a partial minimum $(W_{D}^{*},W_{E}^{*})$ of $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibiting this pattern with positive loss.
@@ 3.5 Theoretical Explanation for Feature Absorption
Feature absorption—where one neuron captures, or ”absorbs”, a specific sub-concept while another responds to remaining related concepts—frequently occurs in SDL training (Chanin et al., 2025). Though prevalently encountered and unwanted, researchers have limited understanding about why feature absorption occurs in SDL training. Here we show hierarchical concept structures naturally introduce realizable activation patterns exhibiting feature absorption and therefore connected with the framework’s partial minima.
Figure 5: Feature absorption emerges from hierarchical concept structure. Left: Ideal SDL features without absorption. Right: hierarchical concept structure exists and only a proportion of the sub-concepts of ”Dog” can activate the SDL feature.
@@
Consider a representation space with a parent concept ”Dog” and four sub-concepts: ”Border Collie”, ”Golden Retriever”, ”Husky”, and ”German Shepherd”. Ideally, SDL learns separate monosemantic neurons for Dog, Cat, Horse, and Elephant, with each dog breed activating only the Dog neuron. However, feature absorption can result in the pattern in Figure 5: one neuron exclusively captures ”Border Collie” (absorbed feature), while another responds to the remaining three breeds collectively (main line interpretation).
@@ 4.1 Anchor Feature Extraction
For real-world datasets where ground-truth features are unknown, we identify semantic subpopulations Luo et al. (2024) and compute their mean representations. Specifically, given a labeled dataset $\mathcal{D}=\{(s_{j},y_{j})\}_{j=1}^{N}$, where $y_{j}\in\{1,\ldots,C\}$, we compute the mean of representations for each class:
@@ 4.2 SDL Loss Function with Feature Anchoring
This feature anchoring technique (Equation 25) reduces the underdetermined nature of SDL training, and is method-agnostic: it applies equally to SAEs, transcoders, crosscoders, and their variants (TopK SAEs, Matryoshka SAEs, etc.) since it only constrains the encoder $W_{E}$ and decoder $W_{D}$ matrices that are common to all SDL architectures. This universality makes feature anchoring a broadly useful technique for improving performance across SDL methods.
@@ 5 Experimental Results
We first evaluate feature anchoring with various SDL methods on the Linear Representation Bench (Section 5.1), a synthetic benchmark with fully accessible ground-truth features that precisely instantiates our theoretical assumptions. Second, we validate feature anchoring on real-world data by training SDL methods on CLIP embeddings of ImageNet-1K. Third, we provide empirical evidence showing that dead neuron resampling helps escape spurious local minima in large language models. We provide comprehensive ablation studies in Appendix G.
@@ 5.1 Results on The Linear Representation Bench
To validate our theoretical predictions, we design the Linear Representation Bench, a benchmark that precisely instantiates Assumptions 2.3 and 2.4 with fully known GT features.
@@
Metrics. Given learned features $W_{p}^{\text{learned}}=W_{E}^{\top}\in\mathbb{R}^{n_{p}\times n_{q}}$ and ground truth features $W_{p}^{\text{true}}\in\mathbb{R}^{n_{p}\times n}$ (columns are unit-norm feature directions), we compute the similarity matrix $\mathbf{S}=|W_{p}^{\text{learned}\top}W_{p}^{\text{true}}|\in\mathbb{R}^{n_{q}\times n}$. For each ground truth feature $i\in[n]$, we find its best match: $s_{i}=\max_{j}S_{ji}$.
@@
  • Maximum Inner Product is the mean of best matches: $\mathbf{M_{IP}}(W_{p}^{learned})=\frac{1}{n}\sum_{i=1}^{n}s_{i}$.
As shown in Table 1, feature anchoring significantly improves feature recovery performance for all SDL methods.
Table 1: Feature recovery on the Linear Representation Bench ($n=1000$, $n_{p}=n_{r}=768$, $n_{q}=16000$, $\mathcal{S}=0.99$). Feature anchoring (FA) consistently improves GT Recovery and Maximum Inner Product (MIP) across all SDL methods.
[…]
ReLU SAE 0.00% 0.205 2787.4
   + Feature Anchoring 0.00% 0.246 1715.1
JumpReLU SAE 0.00% 0.237 1572.3
   + Feature Anchoring 0.00% 0.333 1139.6
TopK SAE 84.90% 0.983 54.1
   + Feature Anchoring 87.63% 0.986 54.1
BatchTopK SAE 84.80% 0.981 65.0
   + Feature Anchoring 89.38% 0.988 60.1
Matryoshka SAE 83.70% 0.982 58.9
   + Feature Anchoring 87.32% 0.985 57.9
Transcoder 23.60% 0.838 50.6
   + Feature Anchoring 25.05% 0.838 49.5
Crosscoder 56.42% 0.940 83.4
   + Feature Anchoring 57.71% 0.941 116.3

### v5

@@ 1 Introduction
Despite this empirical success across diverse applications, SDL methods consistently exhibit persistent failure modes: learned features remain polysemantic (Chanin et al., 2025), "dead neurons" fail to activate on any data samples (Bricken et al., 2023), and "feature absorption" occurs where one neuron captures specific sub-concepts while another responds to the remaining related concepts (Chanin et al., 2025). Practitioners have developed techniques to address these issues, including neuron resampling (Bricken et al., 2023), auxiliary losses (Gao et al., 2024), and careful hyperparameter tuning, yet these fixes remain ad-hoc engineering solutions without principled justification. Critically, these phenomena persist even with careful training on clean data, suggesting they are not mere implementation artifacts but reflect fundamental structural properties of the SDL optimization problem itself.
@@
In this work, we develop a unified theoretical framework that formalizes SDL as a general optimization problem, encompassing various SDL methods (Bussmann et al., 2024; 2025; Tang et al., 2025b; Luo et al., 2023) as special cases. We demonstrate how these diverse methods instantiate our framework through different choices of input-output representation pairs, activation functions, and loss designs. We establish rigorous conditions under which SDL methods provably recover ground-truth interpretable features, characterizing the roles of feature sparsity, latent dimensionality, and activation functions. Through detailed analysis of the optimization landscape, we demonstrate that global minima correspond to correct feature recovery and provide necessary and sufficient conditions for achieving zero loss. We establish the prevalence of spurious partial minima exhibiting un-disentangled polysemanticity, providing novel theoretical explanations for feature absorption (Chanin et al., 2025) and the effectiveness of neuron resampling (Bricken et al., 2023). We design the Linear Representation Bench, a synthetic benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theoretical insights, we propose feature anchoring, a technique applicable to all SDL methods which achieves improved feature recovery by constraining learned features to known anchor directions.
@@ 2 Preliminaries
In this section, we present a unified theoretical framework for Sparse Dictionary Learning (SDL). We begin with the formal definitions of foundational concepts and our assumptions on representations. Then we introduce our framework and how various SDL methods instantiate it.
@@ 2.1 Input Distribution and Model Representation
2.2 Assumptions
Empirical studies in mechanistic interpretability have observed that neural network representations encode meaningful concepts as linear directions, often in superposition (Marks and Tegmark, 2024; Nanda et al., 2023; Jiang et al., 2024; Park et al., 2025). Following Elhage et al. (2022) and Park et al. (2024), we primarily consider model representations $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ satisfying the following assumptions.
Assumption 2.1 (Representation Assumptions).
There exists a feature function $\mathbf{x}:\mathcal{X}\to\mathbb{R}^{n}$ and a feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ such that:
@@ 2.2 Assumptions
We refer to $\mathbf{x}(s)$ as the ground-truth features, where each $x_{i}(s)$ represents the activation level of concept $i$ for input $s$. Researchers typically assume each component $x_{i}$ of $\mathbf{x}$ corresponds to a human-interpretable concept (Bricken et al., 2023; Cunningham et al., 2023; Elhage et al., 2022).
@@ 2.3 General Optimization Framework for SDL
Now we formalize SDL as an optimization problem under the Representation Assumptions (Assumption 2.1). We follow previous approaches Cui et al. (2025); Elhage et al. (2022) to omit the bias terms for mathematical simplicity.
@@ 2.4 Instantiations: Existing SDL Methods
This property holds for ReLU, JumpReLU, Top-$k$, Batch Top-$k$, and their compositions, the primary sparsity mechanisms used in practice.
@@ 3.1 Approximate Loss with Feature Reconstruction
We establish the bound for $S$ satisfying $(1-S)\leq\frac{1}{n}$, the ${\mathcal{L}}_{\text{SDL}}$ can be approximated by per-feature reconstruction.
@@
Under Assumption 2.1, let $C=\sup_{s\sim\mathcal{D}}\|\mathbf{x}_{r}(s)-W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))\|_{2}^{2}<\infty$. Define the approximate loss as:
@@
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}\mid x_{d}(s)>0]$. Then for any $S$ satisfying $(1-S)\leq\frac{1}{n}$:
$\left|\mathcal{L}_{\text{SDL}}-\tilde{\mathcal{L}}_{\text{SDL}}\right|\leq n^{2}C(1-S)^{2}$ (12)
This approximation isolates the contribution of each ground-truth feature, making the optimization landscape amenable to analysis techniques applied subsequently.
@@ 3.2 SDL Optimization is Piecewise Biconvex
Consider the approximate loss $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$. For an activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ (See Definition 3.1), define the corresponding activation pattern region as:
$\Omega_{\mathcal{P}}=\Bigl\{$ $W_{E}\in\mathbb{R}^{n_{q}\times n_{p}}\;\Big|\;\forall i\in[n_{q}],$
$\mathcal{F}_{i}=\bigl\{d\in[n]:\bigl(\sigma(W_{E}w_{p}^{d})\bigr)_{i}>0\bigr\}\Bigr\},$ (13)
Then $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibits bi-convex structure over $\mathbb{R}^{n_{r}\times n_{q}}\times\Omega_{\mathcal{P}}$:
@@
    For any fixed $W_{E}\in\Omega_{\mathcal{P}}$, $W_{D}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{D}$.
@@
    For any fixed $W_{D}\in\mathbb{R}^{n_{r}\times n_{q}}$, $W_{E}\mapsto\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ is convex in $W_{E}$ over $\Omega_{\mathcal{P}}$.
This establishes SDL as a biconvex optimization problem within each activation pattern, bridging mechanistic interpretability with classical biconvex optimization theory, enabling both theoretical analysis of the optimization landscape and implementation of biconvex methods for SDL.
@@ 3.3 Characterizing the Global Minimum of SDL
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}\mid x_{d}(s)>0]$, and let $M=\max_{i\neq j}\langle w_{p}^{i},w_{p}^{j}\rangle$ denote the maximum interference. When $n_{q}\geq n$ and $\sigma$ satisfies $\sigma(z)_{i}\in\{0,z_{i}\}$ for all $i$, the configuration
@@
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})\leq n^{2}M^{2}\sum_{d=1}^{n}M_{d}$ (16)
@@ 3.4 Characterizing Spurious Partial Minima of SDL
Under Assumptions 2.1 with $n\geq 2$ and $n_{q}\geq n$, for any activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ that is realizable, polysemantic, and forms a partition of $[n]$, there exists a partial minimum $(W_{D}^{*},W_{E}^{*})$ of $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibiting this pattern with positive loss.
@@ 3.5 Theoretical Explanation for Feature Absorption
Feature absorption—where one neuron captures, or "absorbs", a specific sub-concept while another responds to remaining related concepts—frequently occurs in SDL training (Chanin et al., 2025). Though prevalently encountered and unwanted, researchers have limited understanding about why feature absorption occurs in SDL training. Here we show hierarchical concept structures naturally introduce realizable activation patterns exhibiting feature absorption and therefore connected with the framework’s partial minima.
@@
Consider a representation space with a parent concept "Dog" and four sub-concepts: "Border Collie", "Golden Retriever", "Husky", and "German Shepherd". Ideally, SDL learns separate monosemantic neurons for Dog, Cat, Horse, and Elephant, with each dog breed activating only the Dog neuron. However, feature absorption can result in the pattern in Figure 5: one neuron exclusively captures "Border Collie" (absorbed feature), while another responds to the remaining three breeds collectively (main line interpretation).
@@
Figure 5: Feature absorption emerges from hierarchical concept structure. Left: Ideal SDL features without absorption. Right: hierarchical concept structure exists and only a proportion of the sub-concepts of "Dog" can activate the SDL feature.
@@ 4.1 Anchor Feature Extraction
For real-world datasets where ground-truth features are unknown, we identify semantic subpopulations Luo et al. (2024b) and compute their mean representations. Specifically, given a labeled dataset $\mathcal{D}=\{(s_{j},y_{j})\}_{j=1}^{N}$, where $y_{j}\in\{1,\ldots,C\}$, we compute the mean of representations for one subpopulation cluster, which is considered as a proxy for ground-truth features:
@@ 4.2 SDL Loss Function with Feature Anchoring
This feature anchoring technique (Equation 26) reduces the underdetermined nature of SDL training, and is method-agnostic: it applies equally to SAEs, transcoders, crosscoders, and their variants (TopK SAEs, Matryoshka SAEs, etc.) since it only constrains the encoder $W_{E}$ and decoder $W_{D}$ matrices that are common to all SDL architectures. We believe this universality makes feature anchoring a broadly useful technique for improving performance across SDL methods.
@@ 5 Experimental Results
We first evaluate feature anchoring with various SDL methods on the Linear Representation Bench (Section 5.1), a synthetic benchmark with fully accessible ground-truth features that precisely instantiates our theoretical assumptions. Second, we validate feature anchoring on real-world data by training SDL methods on CLIP embeddings of ImageNet-1K. Third, we provide empirical evidence showing that dead neuron resampling helps escape spurious local minima in large language models. Ablation studies are in Appendix G.
@@ 5.1 Results on The Linear Representation Bench
To validate our theoretical predictions, we design the Linear Representation Bench, a benchmark that precisely instantiates Assumptions 2.1 with fully known GT features.
@@
Metrics. Given learned features $W_{p}^{\text{learned}}=W_{E}^{\top}\in\mathbb{R}^{n_{p}\times n_{q}}$ and ground truth features $W_{p}^{\text{true}}\in\mathbb{R}^{n_{p}\times n}$ (columns are unit-norm feature directions), we compute the similarity matrix $\mathbf{S}=|W_{p}^{\text{learned}\top}W_{p}^{\text{true}}|\in\mathbb{R}^{n_{q}\times n}$. For each ground truth feature $i\in[n]$, we find its best match score: $s_{i}=\max_{j}S_{ji}$.
@@
  • Maximum Inner Product is the mean of best match scores: $\mathbf{M_{IP}}(W_{p}^{learned})=\frac{1}{n}\sum_{i=1}^{n}s_{i}$.
As shown in Table 1 and Figure 6, feature anchoring significantly improves feature recovery performances.
Table 1: Feature recovery results on the Linear Representation Bench ($n=1000$, $n_{p}=n_{r}=768$, $n_{q}=16000$, $\mathcal{S}=0.99$). Evaluation is performed only on features not used for anchoring. Feature anchoring (FA) consistently improves both GT Recovery ($\mathbf{M_{GT}}$) and Maximum Inner Product ($\mathbf{M_{IP}}$) across all SDL methods. *: For some methods, $\mathbf{M_{GT}}$ is 0% because all best-match similarities fall below the evaluation threshold (0.95). In these cases, $\mathbf{M_{IP}}$, which remains well-aligned with $\mathbf{M_{GT}}$, serves as a more informative indicator of recovery performance. See Figure 6 for a threshold-dependent view of $\mathbf{M_{GT}}$.
Method $\mathbf{M_{GT}}$ $\uparrow$ $\mathbf{M_{IP}}$ $\uparrow$
ReLU SAE 0.00%* 0.205
   + Feature Anchoring 0.00%* 0.246
JumpReLU SAE 0.00%* 0.237
   + Feature Anchoring 0.00%* 0.333
TopK SAE 84.90% 0.983
   + Feature Anchoring 87.63% 0.986
BatchTopK SAE 84.80% 0.981
   + Feature Anchoring 89.38% 0.988
Matryoshka SAE 83.70% 0.982
   + Feature Anchoring 87.32% 0.985
Transcoder 23.60% 0.838
   + Feature Anchoring 25.05% 0.838
Crosscoder 56.42% 0.940
   + Feature Anchoring 57.71% 0.941
Figure 6: $\mathbf{M_{GT}}$ vs. Threshold. Across all SDL methods, feature anchoring consistently improves feature recovery, as measured by $\mathbf{M_{GT}}$ over a range of thresholds.